Write the equation ydx+ (x-y+x-2y)=0 in the linear form and hence find its solution.
1
Expert's answer
2020-12-10T13:50:10-0500

"\displaystyle
y\mathrm{d}x+ (x-y+x-2y)\mathrm{d}y =0 \\

y\mathrm{d}x+ (2x - 3y)\mathrm{d}y =0 \\


y\mathrm{d}x+ (2x - 3y)\mathrm{d}y =0 \\

\frac{\mathrm{d}x}{\mathrm{d}y} = \frac{3y - 2x}{y}\\


\frac{\mathrm{d}x}{\mathrm{d}y} = 3 - \frac{2x}{y}\\



\frac{\mathrm{d}x}{\mathrm{d}y} + \frac{2x}{y} = 3\, \textsf{as a linear form}\\


\textsf{Integrating factor}\,\, (IF) = e^{\int \frac{2}{y}\,\mathrm{d}y} = e^{2\ln{y}} = e^{\ln(y^2)} = y^2\\


\frac{\mathrm{d}x}{\mathrm{d}y} + \frac{2x}{y} = 3\\


xIF = \int 3 IF\, \mathrm{d}y\\

xy^2 = \int 3y^2 \, \mathrm{d}y\\


xy^3 = y^3 + C\\

x = 1 + Cy^{-3}\, \textsf{is the solution to the ODE}"